# Bo Du, Mang Ye

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is widely adopted for downstream fine-tuning of foundation models due to its efficiency and zero additional inference cost. Many real-world applications require foundation models to specialize in multiple tasks simultaneously, motivating the need for efficient multi-task adaptation. While recent approaches integrate LoRA with mixture-of-experts (MoE) to address this, the use of routers prevents parameter mergeability, which increases inference overhead and hinders unified multi-task adaptation, thereby limiting deployment practicality. In this work, we propose ThanoRA, a Task Heterogeneity-Aware Multi-Task Low-Rank Adaptation framework that enables multi-task adaptation while preserving the inference efficiency of LoRA. ThanoRA jointly models task heterogeneity and mitigates subspace interference throughout training. Specifically, motivated by inherent differences in complexity and heterogeneity across tasks, ThanoRA constructs task-specific LoRA subspaces at initialization, enabling fine-grained knowledge injection aligned with task heterogeneity. Furthermore, to prevent task interference and subspace collapse during multi-task training, ThanoRA introduces a subspace-preserving regularization that maintains the independence of task-specific representations. With the synergy of both components, ThanoRA enables efficient and unified multi-task adaptation. Extensive experiments across multimodal and text-only benchmarks under varying multi-task mixtures demonstrate that ThanoRA consistently achieves robust and superior performance over strong baselines without introducing additional inference overhead. Our code is publicly available at: https://github.com/LiangJian24/ThanoRA.
@@ 1 Introduction
With the rapid development of foundation models li2024llava ; LLaVA15_CVPR23 ; wang2024qwen2 ; chen2024internvl ; touvron2023llama ; grattafiori2024llama , their general capabilities across a wide range of tasks have been significantly enhanced. Nevertheless, they often require downstream adaptation to perform well on specific target tasks LoRA_ICLR22 ; huang2025keeping . To address this, a variety of parameter-efficient fine-tuning (PEFT) methods have been proposed to adapt foundation models with minimal computational and storage overhead houlsby2019parameter ; lester2021power ; liu2022few ; li2021prefix ; LoRA_ICLR22 . Among them, LoRA LoRA_ICLR22 and its variants meng2024pissa ; DoRA_ICML24 ; yang2024corda ; hayou2024lora+ are particularly appealing due to their low training cost and the ability to merge adapted parameters into the base model without incurring inference overhead. While prior PEFT research has primarily focused on single-task adaptation, real-world applications often demand that a single model specializes in several specific and complementary capabilities simultaneously li2024mixlora ; liu2024moe . For example, document understanding may require both visual recognition (e.g., OCR) and complex textual reasoning. This motivates the need for multi-task PEFT, where the model acquires multiple skills without sacrificing efficiency and effectiveness.
Recently, a number of studies have combined LoRA with mixture-of-experts (MoE) architectures to achieve multi-task adaptation of foundation models gou2023mixture ; li2024mixlora ; shen2024multimodal ; zhao2025each ; tian2024hydralora ; chen2024llava ; wu2024mixture . These approaches introduce routers to dynamically dispatch different task samples to independent LoRA experts, thereby alleviating task conflicts and thus improving multi-task performance. Although effective to some extent, the additional router structures prevent the adapted parameters from being merged back into the foundation model (as shown in Figure 1(a)), resulting in increased inference latency—an issue that becomes more severe as model sizes continue to grow LoRA_ICLR22 . In contrast, preserving parameter mergeability not only eliminates additional inference overhead but also facilitates the seamless integration of multi-task adaptations into a unified and versatile foundation model.
To this end, this work aims to develop a multi-task adaptation framework without introducing additional non-mergeable structures (as illustrated in Figure 1(b)). Specifically, we focus on addressing the following two key challenges:
Challenge I) How to enable multi-task adaptation without relying on additional structure?
In a unified LoRA framework without additional non-mergeable structure, where different tasks share the same set of trainable parameters without the support of external structures like routers, task conflicts and performance degradation may easily occur. To address this, we seek to explicitly model task heterogeneity to guide parameter subspace allocation and mitigate interference. This requires understanding the distinct representational demands of different tasks, particularly in terms of their complexity and intrinsic task heterogeneity.
Figure 1: Motivation. (a) MoE-based LoRA mitigates task conflicts via routers but compromises parameter mergeability. (b) Our method enables efficient multi-task adaptation through mergeable LoRA with task-specific subspace modeling. (c) Tasks show heterogeneous complexity across layers, motivating the allocation of task-specific subspaces to better match their representational demands.
@@
We first observe that different tasks often exhibit varying degrees of differences in complexity and characteristics across various aspects. For example, as shown in Figure 1(c), some tasks involve simpler visual content but more complex textual information, while others present the opposite pattern. Such heterogeneity indicates that different tasks have distinct demands on the representational capacity at different layers. Consequently, applying a uniform adaptation strategy across tasks may fail to capture these complexity differences, resulting in suboptimal resource allocation. Motivated by this observation, we propose Heterogeneity-Aware Subspace Initialization (HASI). Specifically, task-specific priors are extracted based on instruction-previewed SVD, as similarly employed in yang2024corda . We then estimate task heterogeneity via the spectral entropy of singular values, which serves as a proxy for the subspace complexity required by each task. Based on the relative heterogeneity of each task, we partition the $R$-dimensional subspace spanned by the LoRA modules at each layer. Each task-specific subspace is then initialized with the truncated singular vectors corresponding to its allocated dimension, enabling fine-grained, layer-wise knowledge injection tailored to task complexity.
While HASI effectively constructs independent task-specific subspaces during initialization, shared parameter updates across mixed-task samples may still cause interference between task-specific subspaces, leading to knowledge interference and subspace collapse, ultimately destroying the task-specific structure (see Sec. 4.3.3). Thus, this raises the second key challenge:
@@
Challenge II) How to prevent task subspace collapse and interference during mixed-task training?
To mitigate this, we further propose Subspace-Preserving Regularization (SPR), which introduces cross-task orthogonality regularization during training to distribute task knowledge across independent subspaces, while incorporating a small collaborative rank to preserve shared structure. This suppresses interference and collapse among subspaces, promoting stable and collaborative multi-task adaptation.
Notably, HASI and SPR work synergistically: HASI explicitly preserves task-specific parameter subspaces at initialization, providing a structural foundation that enables SPR to maintain their independence throughout training. Our contributions are summarized as follows:
  • Revealing Task Complexity and Heterogeneity. We reveal that downstream tasks exhibit varying levels of complexity and characteristics, which limits the effectiveness of uniform adaptation and motivates differentiated adaptation within a unified framework.
@@
  • Parameter-Mergeable Multi-Task LoRA Framework. We propose Heterogeneity-Aware Subspace Initialization and Subspace-Preserving Regularization to address the challenges of task heterogeneity modeling and subspace collapse during multi-task adaptation, enabling effective multi-task adaptation while preserving parameter mergeability and inference efficiency.
@@
  • Comprehensive Empirical Validation across Diverse Tasks. We conduct extensive experiments on multimodal and text-only benchmarks under varying multi-task mixtures, consistently demonstrating robust performance improvements and efficiency advantages over strong baselines.
@@ 2 Related Work
Parameter-Efficient Fine-Tuning (PEFT) methods aim to adapt large foundation models with minimal trainable parameters, avoiding the prohibitive cost of full fine-tuning houlsby2019parameter ; lester2021power ; li2021prefix ; liu2022few ; pfeiffer2020adapterfusion ; karimi2021compacter ; wu2024llama ; LoRA_ICLR22 ; LoRASculpt_CVPR25 . Among these, Low-Rank Adaptation (LoRA) LoRA_ICLR22 is widely adopted due to its simplicity and zero additional inference cost, injecting trainable low-rank matrices into existing weight layers without modifying the model architecture. Numerous extensions have been proposed to improve its performance. DoRA DoRA_ICML24 decomposes pretrained weights into magnitude and direction components to improve stability and adaptation capacity. PiSSA meng2024pissa accelerates convergence by initializing low-rank matrices with the principal components of the pretrained weights instead of random noise. Other approaches include dynamic rank allocation zhang2023adalora , vector-based random matrix adaptation kopiczko2023vera , instruction-previewed initialization yang2024corda . These variants improve the capacity and convergence behavior of LoRA, without sacrificing its efficiency in both training and inference.
@@
Multi-task learning (MTL) aims to enhance generalization by jointly training on multiple tasks, enabling shared representations to improve knowledge integration crawshaw2020multi ; zhang2021survey ; ban2024fair ; agiza2024mtlora ; tang2024merging ; huang2023lorahub . Classical approaches include parameter sharing wallingford2022task ; ma2019snr , task-specific structure guo2020learning ; sun2021task , or regularization-based techniques zhang2014regularization ; yang2016trace . With the rise of large foundation models, efficiently adapting them to diverse downstream tasks has become a key focus in recent MTL research. A growing number of studies have proposed efficient solutions for adapting foundation models to multi-task scenarios, mainly by combining Mixture-of-Experts (MoE) architecture with LoRA li2024mixlora ; shen2024multimodal ; zhao2025each ; tian2024hydralora ; chen2024llava ; wu2024mixture ; feng2024mixture ; wang2024malora ; zhu2025remedy . For example, MixLoRA shen2024multimodal mitigates task interference by conditionally constructing task-aware LoRA factors for task-specific routing. HydraLoRA tian2024hydralora introduces an asymmetric MoE to improve multi-task adaptability and parameter efficiency. While these MoE-based methods alleviate task conflicts to some extent, they rely on non-mergeable additional structure like router, which introduce additional inference overhead and compromise the deployment efficiency advantage of LoRA.
@@ 3.1 Preliminary
Instruction-Previewed Initialization. Following previous work yang2024corda , we extract task-specific priors by forwarding a small batch (e.g., 500) of task-specific samples through the pretrained foundation model. For each linear layer, we collect the input activation $X\in\mathbb{R}^{d_{\text{in}}\times BL}$, compute the covariance $C=XX^{T}$, and derive a task-aware Singular Value Decomposition (SVD) as follows:
@@ 3.2 Proposed Method
To address the challenges of task heterogeneity (see Challenge I) and subspace collapse (see Challenge II) in multi-task adaptation, we propose ThanoRA, a Task Heterogeneity-Aware framework for multi-task low-rank adaptation.
@@ 3.2.1 Heterogeneity-Aware Subspace Initialization (HASI)
Real-world tasks often exhibit notable complexity differences across various aspects, leading to divergent representational demands across foundation model layers. For example, as shown in Figure 1(c), in multi-task fine-tuning of MLLM, DocVQA mathew2021docvqa task presents visually complex but textually simple inputs, whereas ScienceQA ScienceQA_NeurIPS22 task features the opposite. This heterogeneity is further reflected in the normalized spectral entropy curves (computed via Eq. (3)), where each task shows distinct complexity patterns across layers. These observations align with prior findings that early layers of MLLMs focus on cross-modal fusion, while later layers predominantly process the fused semantic representations huang2025keeping ; zhang2024cross . To capture such task-layer discrepancies, we construct task-specific low-rank subspaces aligned with layer-wise complexity.
Based on this motivation, to capture task-specific structural priors, we first obtain a task-aware representation matrix $\widehat{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ using instruction-previewed SVD, as defined in Eq. (2). This captures the dominant semantic directions associated with the task-specific response to the pretrained model. Then, to quantify the complexity of the representation required by each task at a given layer, we compute the spectral entropy of the singular values $\bm{\sigma}\in\mathbb{R}^{r_{\text{total}}}$ of $\widehat{W}$ as:
@@
where $r_{\text{total}}$ denotes the rank budget for truncating the SVD of $\widehat{W}$. This reflects the effective rank and expressiveness of the layer-specific subspace required by each task.
Under a fixed total rank budget $r_{\text{total}}$, we allocate a portion $r_{t}^{\ell}$ of the LoRA rank to each task $t$ at layer $\ell$ proportionally to the normalized entropy across tasks:
@@
The remaining task-specific rank budget $r_{\text{task}}=r_{\text{total}}-r_{\text{coop}}$ is then distributed by discretizing the soft allocations:
@@
where $r_{\min}$ ensures each task receives a minimum allocation to avoid rank collapse.
@@
Having obtained the task-specific components $\mathbf{B}_{t}^{\ell}$ and $\mathbf{A}_{t}^{\ell}$ for each task $t$, we define the corresponding low-rank parameter subspace at layer $\ell$ as
$\mathcal{S}_{t}^{\ell}=\mathrm{Span}(\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}).$ (7)
To construct a unified LoRA while explicitly preserving parameters corresponding to each task, we concatenate all task-specific matrices in a blockwise manner. The proposition below shows that this composition faithfully preserves the union of task-specific subspaces.
Proposition 3.1 (Subspace Equivalence via Blockwise LoRA Composition).
Let $\mathcal{S}_{t}^{\ell}=\mathrm{Span}(\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell})$ denote the task-specific low-rank subspace for task $t$ at layer $\ell$, and let $\mathcal{S}_{\text{total}}^{\ell}=\mathrm{Span}(\mathbf{B}^{\ell}\mathbf{A}^{\ell})$ denote the overall subspace constructed via blockwise LoRA initialization, where
$\mathbf{B}^{\ell}=[\mathbf{B}_{1}^{\ell},\dots,\mathbf{B}_{T}^{\ell}],\quad\mathbf{A}^{\ell}=\left[\mathbf{A}_{1}^{\ell};\cdots;\mathbf{A}_{T}^{\ell}\right].$ (8)
Then the composed subspace satisfies:
$\mathcal{S}_{\text{total}}^{\ell}=\sum_{t=1}^{T}\mathcal{S}_{t}^{\ell}.$ (9)
Proof. By the rule of block matrix multiplication, we have:
$\mathbf{B}^{\ell}\mathbf{A}^{\ell}=\left[\mathbf{B}_{1}^{\ell},\cdots,\mathbf{B}_{T}^{\ell}\right]\cdot\begin{bmatrix}\mathbf{A}^{\ell}_{1}\\ \mathbf{A}^{\ell}_{2}\\ \vdots\\ \mathbf{A}^{\ell}_{T}\end{bmatrix}=\sum_{t=1}^{T}\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}.$ (10)
Each term $\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}$ lies within its respective subspace $\mathcal{S}_{t}^{\ell}$, and thus the span of the sum equals the sum of spans:
$\mathrm{Span}(\mathbf{B}^{\ell}\mathbf{A}^{\ell})=\sum_{t=1}^{T}\mathrm{Span}(\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}).$ (11)
Hence, the LoRA parameter space constructed by direct concatenation of task-specific components faithfully represents the union of the original task-specific subspaces. ∎
We then concatenate all task-specific LoRA components along with the cooperative block:
$\mathbf{B}^{\ell}$ $=\left[\mathbf{B}_{1}^{\ell},\cdots,\mathbf{B}_{T}^{\ell},\mathbf{B}_{\text{coop}}^{\ell}\right],$ (12)
$\mathbf{A}^{\ell}$ $=\left[\mathbf{A}_{1}^{\ell};\cdots;\mathbf{A}_{T}^{\ell};\mathbf{A}_{\text{coop}}^{\ell}\right],$ (13)

### v2

# ThanoRA: Task Heterogeneity-Aware Multi-Task Low-Rank Adaptation

@@ Abstract
Low-Rank Adaptation (LoRA) is widely adopted for downstream fine-tuning of foundation models due to its efficiency and zero additional inference cost. Many real-world applications require foundation models to specialize in several specific tasks simultaneously, motivating the need for efficient multi-task downstream adaptation. To address this need, existing studies have primarily explored two directions: Model Merging with LoRA, which shows advantages in training-free scenarios but still lags behind multi-task training in overall performance; and MoE-based LoRA approaches, which improve multi-task learning performance but introduce routers that hinder the mergeability of LoRA parameters and incur considerable inference overhead, thereby limiting real-world deployment practicality. To this end, we propose ThanoRA, a Task Heterogeneity-Aware Multi-Task Low-Rank Adaptation framework that enables effective, efficient and unified multi-task downstream adaptation without introducing additional structure. ThanoRA performs multi-task learning by tailoring subspace allocation at initialization and enforcing diversity preservation throughout training: it allocates varying dimensions to construct task-specific low-rank subspaces driven by inter-task heterogeneity, enabling fine-grained knowledge injection, while diversity-preserving regularization mitigates task interference and subspace collapse, thereby fully exploiting the low-rank capacity. Extensive experiments across multimodal and text-only benchmarks under varying multi-task mixtures demonstrate that ThanoRA consistently outperforms strong baselines, surpassing even separate task-specific fine-tuning, while introducing no additional structures or inference overhead. Our code will be publicly available at: https://github.com/LiangJian24/ThanoRA.
@@ 1 Introduction
With the rapid development of foundation models Bai et al. (2025); Wang et al. (2025); Grattafiori et al. (2024); Li et al. (2024a); Liu et al. (2023), their general capabilities across a wide range of tasks have been significantly enhanced. Nevertheless, they often require downstream adaptation to perform well on specific target tasks Hu et al. (2022); Huang et al. (2025). To address this, a variety of parameter-efficient fine-tuning (PEFT) methods have been proposed to adapt foundation models with minimal computational and storage overhead Houlsby et al. (2019); Lester et al. (2021); Liu et al. (2022); Li & Liang (2021); Hu et al. (2022). Among them, LoRA Hu et al. (2022) and its variants Meng et al. (2024); Liu et al. (2024b); Yang et al. (2024); Hayou et al. (2024) are particularly appealing due to their low training cost and the ability to merge adapted parameters into the base model without incurring inference overhead. Although PEFT methods such as LoRA have shown competitive performance, recent studies suggest that they may encounter challenges on heterogeneous corpora involving multiple tasks or domains, where task conflicts and interference can hinder adaptation Tian et al. (2024). Meanwhile, real-world applications often require a single model to acquire several targeted capabilities simultaneously Li et al. (2024b); Liu et al. (2024a), motivating the need for efficient multi-task downstream adaptation.
Existing efforts toward multi-task efficient adaptation have primarily explored two directions. The first is Model Merging with LoRA (see fig. 1.a), where individually fine-tuned LoRA modules from different tasks are directly merged to form a multi-task model Stoica et al. (2025); Zeng et al. (2025); Zheng et al. (2025). This approach is attractive due to its training-free nature in combining arbitrary sets of tasks, but the lack of a shared optimization process limits its ability to resolve inter-task conflicts, resulting in inferior performance to multi-task training. The second line of research combines LoRA with Mixture-of-Experts (MoE) (see fig. 1.b) architectures Li et al. (2024b); Shen et al. (2024); Zhao et al. (2025); Tian et al. (2024); Chen et al. (2024); Wu et al. (2024b). These methods employ routers to dynamically dispatch inputs to different LoRA experts, thereby alleviating conflicts and improving multi-task performance. Nevertheless, the reliance on routers prevents the adapted parameters from being merged back into the base model, leading to considerable inference overhead and extra storage requirements, thereby hindering real-world deployment.
To this end, we aim to develop an efficient multi-task low-rank adaptation framework better aligned with real-world deployment, enabling large foundation models to master the combination of targeted tasks without incurring additional inference or storage overhead. This raises the following challenge:
Challenge I) How to effectively inject multi-task knowledge into the low-rank subspace?
Conducting multi-task training in a unified LoRA framework without additional non-mergeable structures requires tasks to share the same trainable parameter subspace, often leading to conflicts and performance degradation. Thus, the key challenge is to compress multi-task knowledge into a shared low-rank space. To address this, we first note the following fact:
Tasks exhibit heterogeneous representational demands across layers of large foundation models.
@@
Prior studies have shown that LoRA modules exhibit varying representational requirements across layers and modules, necessitating different rank configurations Zhang et al. (2023). In the context of multi-task learning, we further extend this conclusion by observing that different tasks present heterogeneous representational demands across different layers. As illustrated in fig. 1.d, consider a concrete example: some tasks involve relatively simple visual content but require complex textual reasoning, whereas others display the opposite pattern. This heterogeneity results in divergent representational capacities required by the foundation model at different layers, as reflected in the accompanying bar chart. While this example illustrates visual–textual heterogeneity, real tasks exhibit multidimensional variations that further accentuate representational differences. Consequently, applying a uniform adaptation strategy in multi-task training may fail to capture these complexity differences, leading to suboptimal resource allocation. This motivates us to allocate low-rank resources in a task-aware manner, aligning the representational capacity at each layer with the heterogeneous demands of different tasks. To this end, we propose Heterogeneity-Aware Subspace Initialization (HASI). Specifically, task-specific priors are extracted based on instruction-previewed SVD, as similarly employed in Yang et al. (2024). We then estimate task heterogeneity via the spectral entropy of singular values, which serves as a proxy for the subspace complexity required by each task. Based on the relative heterogeneity of each task, we partition the $R$-dimensional subspace spanned by the LoRA modules at each layer. Each task-specific subspace is then initialized with the truncated singular vectors corresponding to its allocated dimension, enabling fine-grained, layer-wise knowledge injection tailored to task complexity.
Figure 1: Comparison and Motivation. (a) LoRA Merging acquires multi-task knowledge without training but remains suboptimal compared to multi-task training. (b) MoE-based LoRA mitigates task conflicts via routers but compromises parameter mergeability. (c) Our method enables efficient multi-task adaptation through mergeable LoRA with task-specific subspace modeling. (d) Tasks show heterogeneous complexity across layers, motivating the allocation of task-specific subspaces to better match their representational demands.
While HASI enables a proper allocation of low-rank subspaces and task-aware knowledge injection at initialization, during multi-task training the parameter subspaces of different tasks still share gradient update signals. As a result, the initially disentangled task-specific subspaces may gradually lose their diversity, leading to subspace collapse (see Sec. 3.2.2 for definition). Thus, this raises the second key challenge:
@@
Challenge II) How to mitigate low-rank subspace collapse during multi-task training?
To mitigate this, we further propose Subspace-Preserving Regularization (SPR), which enforces orthogonality among different task-specific subspaces during training while retaining a partially shared collaborative subspace for knowledge sharing. This orthogonality constraint enhances the expressiveness of the low-rank subspaces initialized by HASI and effectively alleviates the problem of subspace collapse (refer to Sec. 4.3.3 for experimental evidence).
Notably, HASI and SPR work synergistically: HASI creates heterogeneous subspaces for task-aware knowledge injection at initialization, while SPR preserves their diversity throughout training, jointly enabling effective multi-task adaptation within a unified, mergeable LoRA framework. Our contributions are summarized as follows:
  • Revealing and Modeling Heterogeneous Representational Demands Across Layers. We reveal that downstream tasks vary in complexity and characteristics across layers, motivating layer-wise heterogeneity modeling within a unified framework.
@@
  • Effective and Efficient Multi-Task Low-Rank Adaptation Framework. We propose ThanoRA to address the challenges of task heterogeneity modeling and subspace collapse during multi-task training, enabling effective multi-task adaptation while preserving parameter mergeability and inference efficiency, supplementing the shortcomings of merging- and MoE-based methods.
@@
  • Comprehensive Empirical Validation. We conduct extensive experiments on multimodal and text-only benchmarks, consistently demonstrating robust performance improvements and efficiency advantages over strong baselines, surpassing even separate task-specific fine-tuning.
@@ 2 Related Work
Recent advances in large foundation models have driven rapid progress Wang et al. (2025); Yang et al. (2025); Rong et al. (2025); Ye et al. (2025); Fang et al. (2025). However, with their ever-increasing scale, full fine-tuning is becoming prohibitively costly in computation and storage. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by efficiently adapting large models with minimal trainable parameters Houlsby et al. (2019); Lester et al. (2021); Li & Liang (2021); Liu et al. (2022); Pfeiffer et al. (2020); Karimi Mahabadi et al. (2021); Wu et al. (2024a); Hu et al. (2022); Liang et al. (2025). Among these, Low-Rank Adaptation (LoRA) Hu et al. (2022) is widely adopted due to its simplicity and zero additional inference cost, injecting trainable low-rank matrices into existing weight layers without modifying the model architecture. Numerous extensions have been proposed to improve its performance. DoRA Liu et al. (2024b) decomposes pretrained weights into magnitude and direction components to improve stability and adaptation capacity. PiSSA Meng et al. (2024) accelerates convergence by initializing low-rank matrices with the principal components of the pretrained weights instead of random noise. Other approaches include dynamic rank allocation Zhang et al. (2023), vector-based random matrix adaptation Kopiczko et al. (2023), instruction-previewed initialization Yang et al. (2024). These variants improve the capacity and convergence behavior of LoRA, without sacrificing its efficiency in both training and inference.
@@
Multi-task learning (MTL) aims to enhance generalization by jointly training on multiple tasks, enabling shared representations to improve knowledge integration Crawshaw (2020); Zhang & Yang (2021); Ban & Ji (2024); Agiza et al. (2024); Huang et al. (2023). Classical approaches include parameter sharing Wallingford et al. (2022); Ma et al. (2019), task-specific structure Guo et al. (2020); Sun et al. (2021), or regularization-based techniques Zhang & Yeung (2014); Yang & Hospedales (2016). With the rise of large foundation models, efficiently adapting them to diverse downstream tasks has become a central focus in recent MTL research. Recent studies on parameter-efficient multi-task adaptation can be broadly divided into two categories: (i) Model Merging with LoRA. These methods directly merge LoRA modules trained on individual tasks Stoica et al. (2025); Zeng et al. (2025); Zheng et al. (2025). While attractive for training-free scenarios, they generally lag behind multi-task training due to the absence of explicit inter-task optimization. (ii) MoE-based LoRA. A growing number of works combine LoRA with Mixture-of-Experts (MoE) architectures Li et al. (2024b); Shen et al. (2024); Zhao et al. (2025); Tian et al. (2024); Chen et al. (2024); Wu et al. (2024b); Feng et al. (2024); Wang et al. (2024); Zhu et al. (2025). For example, MixLoRA Shen et al. (2024) mitigates task interference by conditionally constructing task-aware LoRA factors for task-specific routing, while HydraLoRA Tian et al. (2024) introduces asymmetric experts to improve multi-task adaptability. However, these approaches rely on non-mergeable routing structures, incurring considerable inference and additional storage overhead, thereby undermining the deployment efficiency of LoRA. While LoRA model merging sacrifices overall performance and MoE-based LoRA designs compromise mergeability and efficiency, neither fully addresses the need for effective and practical multi-task adaptation.
@@ 3.1 Preliminary
Instruction-Previewed Initialization. Following previous work Yang et al. (2024), we extract task-specific priors by forwarding a small batch (e.g., 500) of task-specific samples through the pretrained foundation model. For each linear layer, we collect the input activation $X\in\mathbb{R}^{d_{\text{in}}\times BL}$, compute the covariance $C=XX^{T}$, and derive a task-aware Singular Value Decomposition (SVD) as follows:
@@ 3.2 Proposed Method
To address the challenges of incorporating heterogeneous task knowledge into the low-rank subspace (Challenge I) and mitigating subspace collapse (Challenge II) in multi-task low-rank adaptation, we propose ThanoRA, a Task Heterogeneity-Aware framework that enables effective and efficient multi-task adaptation.
Eq. (5)
@@ 3.2.1 Heterogeneity-Aware Subspace Initialization (HASI)
Real-world tasks often exhibit heterogeneity across various aspects, leading to divergent representational demands across layers of large foundation model. For example, as shown in fig. 1(c), in multi-task fine-tuning of MLLM, DocVQA Mathew et al. (2021) task presents visually complex but textually simple inputs, whereas ScienceQA Lu et al. (2022) task features the opposite. This heterogeneity is further reflected in the normalized spectral entropy curves (computed via eq. 3), where each task shows distinct complexity patterns across layers. These observations align with prior findings that early layers of MLLMs focus on cross-modal fusion, while later layers predominantly process the fused semantic representations Huang et al. (2025); Zhang et al. (2024). Notably, this example reflects only a single aspect of heterogeneity, while real-world tasks involve richer and multidimensional forms of variation. Such observations highlight the necessity of explicitly modeling heterogeneity in multi-task adaptation.
To capture such discrepancies, we construct task-specific low-rank subspaces aligned with layer-wise representational heterogeneity. We first obtain a task-aware representation matrix $\widehat{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ using instruction-previewed SVD, as defined in eq. 2. This captures the dominant semantic directions associated with the task-specific response to the pretrained model. Then, to quantify the complexity of the representation required by each task at a given layer, we compute the spectral entropy of the singular values $\bm{\sigma}\in\mathbb{R}^{r_{\text{total}}}$ of $\widehat{W_{t}}$ for task $t$ as:
@@
where $r_{\text{total}}$ denotes the total dimension budget for truncating the SVD of $\widehat{W_{t}}$. This quantifies the representational complexity of the subspace at each layer for a given task.
Furthermore, we compress multi-task knowledge into a shared low-rank space, which in LoRA is naturally constrained by the upper bound $r_{\text{total}}$. Under this fixed total dimension budget $r_{\text{total}}$, we allocate a portion $r_{t}^{\ell}$ of the LoRA rank to each task $t$ at layer $\ell$ proportionally to the normalized entropy across tasks:
@@
The remaining task-specific dimension budget $r_{\text{task}}=r_{\text{total}}-r_{\text{coop}}$ is then distributed by discretizing the soft allocations:
@@
where $r_{\min}$ ensures each task receives a minimum allocation to prevent vanishing ranks for certain tasks.
@@
Having obtained the task-specific components $\mathbf{B}_{t}^{\ell}$ and $\mathbf{A}_{t}^{\ell}$ for each task $t$, we define the corresponding low-rank initialization subspace at layer $\ell$ as $\mathbf{W}_{t}^{\ell}\triangleq\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}.$ We aim to fuse the initialization subspaces of different tasks. To this end, we realize the fusion via the following blockwise concatenation:
$\mathbf{B}^{\ell}=\left[\mathbf{B}_{1}^{\ell},\cdots,\mathbf{B}_{T}^{\ell}\right],\mathbf{A}^{\ell}=\left[\mathbf{A}_{1}^{\ell};\cdots;\mathbf{A}_{T}^{\ell}\right],$ (7)
We demonstrate in the proposition below that this blockwise concatenation is effect-equivalent to the additive composition of the task-specific low-rank components.
Proposition 3.1 (Blockwise composition).
Let $\mathbf{B}_{t}^{\ell}\in\mathbb{R}^{d_{\mathrm{out}}\times r_{t}}$ and $\mathbf{A}_{t}^{\ell}\in\mathbb{R}^{r_{t}\times d_{\mathrm{in}}}$ for tasks $t=1,\dots,T$, and define $\mathbf{B}^{\ell}=[\mathbf{B}_{1}^{\ell},\dots,\mathbf{B}_{T}^{\ell}]\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $\mathbf{A}^{\ell}=[\mathbf{A}_{1}^{\ell};\dots;\mathbf{A}_{T}^{\ell}]\in\mathbb{R}^{r\times d_{\mathrm{in}}}$ with $r=\sum_{t}r_{t}$. Then, for any input $\mathbf{X}\in\mathbb{R}^{d_{\mathrm{in}}\times n}$,
$\mathbf{B}^{\ell}\mathbf{A}^{\ell}\mathbf{X}\;=\;\sum_{t=1}^{T}\mathbf{B}_{t}^{\ell}\mathbf{A}_{t}^{\ell}\mathbf{X}.$ (8)
Proof.
See Appendix C.1. ∎
Thus, we fuse task-specific LoRA initialization subspaces from different tasks with heterogeneous ranks via blockwise concatenation, enabling knowledge injection into low-rank subspaces. Meanwhile, this property preserves the initialization parameter subspace of each task explicitly, facilitating the regularization of subsequent multi-task training. We then concatenate all task-specific LoRA components along with the cooperative block:
$\mathbf{B}^{\ell}=\left[\mathbf{B}_{1}^{\ell},\cdots,\mathbf{B}_{T}^{\ell},\mathbf{B}_{\text{coop}}^{\ell}\right],\mathbf{A}^{\ell}=\left[\mathbf{A}_{1}^{\ell};\cdots;\mathbf{A}_{T}^{\ell};\mathbf{A}_{\text{coop}}^{\ell}\right],$ (9)